calculate the total no. of electrons in 180 ml of water

Grade:12th pass

1 Answers

Arun
5 years ago
180 mL water X 1 g/mL = 180 g water 

moles H2O = 180 g / 18 g/mol = 10 mole H2O 

Number of H2O molecules = 10 mol X 6.02X10^23 molecules/mole= 6.02X10^24 molecules 

Total electrons in 1 H2O molecule = 10 electrons/molecule 

So, 6.02 X 10^24 molecules X 10 electrons/molecule = 6.02 X 10^25 electrons